OpenStudy (anonymous):

....need help for this two problems of differentiation of algebraic functions. 1. y= -3/4x-x^2 2. y =1/4 (2x-3)^-10 please help me solve thisss

OpenStudy (yanasidlinskiy):

\(\Huge\bf \color{orange}{Welcome~to~OpenStudy!!}\hspace{-310pt}\color{cyan}{Welcome~to~OpenStudy!!}\hspace{-307.1pt}\color{midnightblue}{Welcome~to~\color{purple}{Open}}\color{blue}{Study!!}\)

OpenStudy (anonymous):

plwease help me solve

OpenStudy (anonymous):

@matricked

OpenStudy (anonymous):

@cody_123

OpenStudy (anonymous):

\[1.y=\frac{ -3x }{ 4 }-x^2\]

OpenStudy (anonymous):

is this the question

OpenStudy (anonymous):

yes that is

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ -3 }{ 4 }-2x\]

OpenStudy (anonymous):

that the first should i do? what rule will use on that?

OpenStudy (anonymous):

just that d/dx x^n=nx^(n-1)

OpenStudy (anonymous):

ah. okaythe exponent rule? what next in that

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ah okay what next should i do

OpenStudy (anonymous):

\[\frac{ 1 }{ 4(2x-3)^{10} }\]

OpenStudy (anonymous):

is this is the question?

OpenStudy (anonymous):

yes. does the other one is already done?

OpenStudy (anonymous):

wait. the (2x-3)^-10is multiply by both of the fraction

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }*(2x-3)^{-10}\]

OpenStudy (anonymous):

is this question?

OpenStudy (anonymous):

yes that is

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }\frac{ d }{ dx }(2x-3)^{-10}\] \[\frac{ 1 }{ 4 }*(-10)*(2x-3)^{-11}\frac{ d }{ dx }(2x-3)\]

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 }*(-10)*(2x-3)^{-11}*(2)\]

OpenStudy (anonymous):

i will simply substitute?

OpenStudy (anonymous):

what substitute?

OpenStudy (anonymous):

the whole number or its the final answer

OpenStudy (anonymous):

it is the final answer

OpenStudy (anonymous):

ah okay thankyou so much for helping me. god bless you

OpenStudy (anonymous):

thanks